Two point charges + 8q and – q are located at x = 0 and x = L respectively. The location of a point on the x axis at which the net electric field due to these two point chages is zero is
 
(a) L/4
(b)  2L
(c) 4 L
(d) 8 L

Grade:6

1 Answers

Navjyot Kalra
9 years ago

(b) –K2q/(x-L)2 + K8q/x2 = 0 ? 1/ (x - L)2 = 4/ x2

or 1/x – L = 2/x ? x = 2 x – 2 L or x = 2 L